OpenStudy (anonymous):

A machinist creates a washer by drilling a hole through the center of a circular piece of metal. If the piece of metal has a radius of x + 10 and the hole has a radius of x + 6, what is the area of the washer? A.x^2-8pie x-64 pie B.x^2=8pie x 64pie C.8pie x-64 pie D.8pie x +64pie

OpenStudy (campbell_st):

Area of Washer = Area of Metal - Area of the hole \[A = \pi (x+10)^2 - \pi(x+6)^2\] this will expand to \[A = \pi(x^2 + 20x + 100 - x^2 - 12x -36)\] collecting like terms \[A=\pi(8x + + 64)\] or \[A=8\pi(x + 8)\]

OpenStudy (anonymous):

wait im lost

OpenStudy (campbell_st):

this uses the area of a circle \[A = \pi r^2\] circular metal disk has a radius x + 10 Hole radius x + 6 washer area = area of the metal disk - area of the hole... just substitute and simplify